car A can cover a distance of X Km in 1 litre petrol & car B can cover a distance of (X+5)Km in 1 Litre petrol.if car A consumes 4 Litres Of petrol more in a journey of 400 kms than car B. find the petrol consume by car B.

Grade:12

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Petrol is consumed on the basis of distance covered. First find the distance covered by both the cars.

Use distance= speed*time

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

Sudheesh Singanamalla
13 years ago

distance travelled by car A = distance travelled by car B = 400Km

petrol required to travel X km = 1L for car A

pertrol required to travel X+5 km = 1L for car b

1L = X+5 km

1km = 1/ X+5

petrol required for B to travel 400km = 400/X+5

petrol required for A to travel 400km = 400/X

given that

400/X - 4 = 400/X+5

(400-4X)/X = 400/ X+5

(400-4X)(X+5) = 400 X -----------------------BY CROSS MULTIPLICATION

now solve the equation, you will get two answers .

the negative answer is omitted and the positive answer is taken as the petrol consumed by car A

add 4 to it and you will get the petrol consumed by car B

 

 

PLEASE APPROVE !